# Subsystem Information Capacity in Random Circuits and Hamiltonian Dynamics

### v2

@@ II.3.3 Relation with OTOC
The SIC proposed in this work has several advantages over the commonly used probe OTOC for diagnosing information scrambling. From experimental perspectives, SIC only requires one call on the evolution operator $U$, thereby eliminating the need for an extra invocation on reversed time evolution $U^{\dagger}$, which is typically required for measuring OTOCs Mi et al. (2021). While OTOCs of local operators are frequently utilized to investigate when and how local information could reach distant locations, the SIC provides a more comprehensive insight into the global structure of information propagation, evolution, and distribution. Notably, this research is the first to unveil the steady-state information structures across a variety of dynamical phases with the lens of SIC.
II.3.4 Relation with tripartite mutual information
Tripartite mutual information has been proposed to explore the intrinsic information scrambling capacity of unitary evolutions Hosur et al. (2016). In settings for TMI, all system qubits are entangled with external ancilla qubits, as already discussed in Sec. II.3.3.
Indeed, if $\rho_{\bar{E}}$ defined in Eq. (9) is $I$ instead of a pure state $|\psi_{\bar{E}}\rangle\langle\psi_{\bar{E}}|$, our setup can be reduced to the case in Hosur et al. (2016), where TMI is given as
@@ II.3.4 Relation with tripartite mutual information
Figure 2: Late-time subsystem information capacity with one-to-one encoding scheme as a function of the subsystem ratio $L_{A}/L$, highlighting the qualitative differences in behavior across various dynamical phases. Each panel presents the x-axis as $L_{A}/L$ and the y-axis as the average late-time SIC, with the initial value 2. (a) Extended phase of non-interacting fermion, with some integrable models following similar trends. (b) Anderson localization phase of non-interacting fermion. Anderson localization phase induced by random Clifford Floquet circuit also follows similar curves. (c) Topological phase with edge states with the entangled qubit at the spatial boundary. (d) Chaotic phases including random circuits and thermal phases of non-integrable interacting systems. Notably, some non-thermal phases of integrable interacting systems also display comparable behavior. (e) Noise-resilience phase in monitored random circuits. (f) Many-body localization phases.
II.3.5 Relation with Holevo information
Another closely related information metric is Holevo information Yuan et al. (2022). The quantity provides a lower bound on the classical information transmission capacity for a quantum channel, analogous to how quantum coherent information bounds the quantum information transmission capacity. The definition of Holevo information, in terms of the effective subsystem channel introduced in this work, is
@@ II.3.5 Relation with Holevo information
Accordingly, under the one-to-one encoding scheme, the first term of Holevo information is reduced to $S_{A}$ in the language of our main settings with reference qubits $R$. The second term of Holevo information is straightforward as the steady-state entanglement for subsystem $A$ averaged over different initial product states. According to the statistical model mapping introduced in Sec. III.1, the second term differs from $S_{A}$ term in our main settings by the free bottom boundary condition for $E$. For comparison, the bottom boundary condition for $E$ for $S_{A}$ and $S_{AR}$ are $\mathbb{I}$ and $\mathbb{C}$, respectively (Notations on the effective statistical model are introduced in Sec. III.1). Holevo information ranges from 0 to 1, with the two limits implying complete information loss and complete information protection, respectively.
@@
 Phase  Entanglement growth  Steady-state entanglement  Steady-state information distribution
Chaotic phase (Sec. III.2.1, IV.3.1) Linear growth Volume law (Haar value) (d) in Fig. 2
Many-body localization (Sec. IV.3.1) Logarithmic growth Volume law (smaller than Haar value) (f) in Fig. 2
Anderson localization (Sec. III.2.2, IV.2) No growth Area law (b) in Fig. 2
Noise-resilient phase (Sec. III.2.1) Linear growth Volume law (smaller than Haar value) (e) in Fig. 2
Free fermion extended phase (Sec. IV.2) Linear growth Volume law (a) in Fig. 2
Interacting integrable systems (Sec. IV.3.2) Linear growth Volume law (d) in Fig. 2
Topological edge states (Sec. IV.4) Linear growth Volume law (smaller than Haar value) (c) in Fig. 2
SWAP circuits (Sec. IV.2) No growth Area law (a) in Fig. 2
CNOT circuits (Sec. III.2.3) No growth Area law (possible long-range entanglement) Fig. 12
II.4 Main results
Following the presentation of our general frameworks, we now summarize the principal findings of our investigation. The late-time SIC curve with varying subsystem ratio $L_{A}/L$ is one of the key focuses of our interest as it reflects the stable spatial distribution of initially localized information. With one-to-one encoding scheme, the late-time SIC curve can show qualitatively distinct behaviors for different types of dynamics as concisely encapsulated in Fig. 2. In other words, by treating evolution dynamics as a ’black box’, our information probe offers a straightforward and insightful means to discern the intrinsic properties of the dynamical phase. This metric paves the way for a deeper comprehension of quantum dynamics and quantum information processing.
We further summarize the results for various quantum many-body phases investigated in this work in Table 1 where entanglement growth and steady-state entanglement for quench dynamics from product states are presented with the steady-state information distribution $MI(x,t\rightarrow\infty)$ under one-to-one encoding scheme. We find that information dynamics and steady-state information distribution exhibit more patterns than entanglement metrics and are better at differentiating dynamical phases. In general, localized eigenmodes for evolution operators tend to suppress steady-state entanglement and can induce effective information trapping. For example, with all eigenstates localized in the one-dimensional AA model, the final information is trapped in the proximity of the initial information encoding position. Additionally, for a system with mixed localized eigenstates and extended eigenstates, a portion of information could be trapped around these localized modes as shown in the case of topological edge states.
In the subsequent sections, we present our findings for the random circuit systems and Hamiltonian quench dynamics. The numerical data were obtained using state-of-the-art computational tools, including the Julia package QuantumClifford.jl for large-scale Clifford circuit simulations, the Python package TensorCircuit Zhang et al. (2023) for scalable free fermion simulations, and the Python package QuSpin Weinberg and Bukov (2017, 2019) for simulations involving interacting fermions.
@@ II.4 Main results
III Results for random circuit dynamics
Random quantum circuits Fisher et al. (2023) emerge as a promising and powerful platform for exploring a variety of non-equilibrium phenomena, particularly from information perspectives including entanglement growth Nahum et al. (2017, 2018a) and information scrambling characterized by OTOC and other metrics Nahum et al. (2018b); von Keyserlingk et al. (2018); Rakovszky et al. (2018); Khemani et al. (2018); Zhuang et al. (2023). Recently, the noise-resilience phase in monitored random circuits that can protect information from mid-circuit measurements also attracted a lot of interest Li et al. (2018, 2019); Skinner et al. (2019a); Ippoliti and Khemani (2021); Lu and Grover (2021); Ippoliti et al. (2022); Gullans and Huse (2020); Choi et al. (2020); Chan et al. (2019); Szyniszewski et al. (2019); Fan et al. (2021); Li and Fisher (2021); Hoke et al. (2023); Turkeshi et al. (2020); Sierant et al. (2022). We remark that mid-circuit measurement is the only ingredient we consider in this work that leads to open system evolution and information leakage, i.e. $I(|A|=L:R)(t)<2$. It is an interesting future direction to explore possibilities and properties of other gadgets leading to information leakage in information dynamics, which include quantum noises Weinstein et al. (2022); Liu et al. (2023a, 2024a, 2024b); Lovas et al. (2023); Turkeshi and Sierant (2024) or construction with explicit ancillary environments Weinstein et al. (2023).
III.1 Analytical framework
In this subsection, we provide a concise review of the analytical method that establishes the correspondence between random quantum circuits and effective classical spin statistical models, with further details elaborated in Appendix C. Within this framework, the entanglement entropy is interpreted as a so-called entanglement membrane in space-time Jonay et al. (2018); Zhou and Nahum (2020); Sierant et al. (2023), i.e. minimal free energy domain walls between different permutation valued spin in one dimension case. For an in-depth discussion of the mapping and the field theory, we direct the reader to Skinner et al. (2019b); Zhou and Nahum (2019); Bao et al. (2020); Jian et al. (2020); Weinstein et al. (2022); Liu et al. (2024a).
The quantum circuit under investigation is composed of random two-qudit unitary gates arranged in a brick-wall layout as shown in Fig 1(a). In this work, we adopt the convention by taking the unit time $\Delta t$ = 1 for an even-odd layer of two-qudit gates. Note that the convention gives the lightcone speed $v=2$, which may be different from some literature where $\Delta t=1$ refers to one even or one odd layer. After $T$ layers of gates, the output density matrix $\rho$ is
@@ III.1 Analytical framework
where $S_{r}$ is the permutation group of degree $r$, $d$ is the local Hilbert space dimension of qudit, and $\text{Wg}_{d^{2}}^{(r)}$ is the Weingarten function Zhou and Nahum (2019); Collins and Śniady (2006). By independently averaging over all two-qudit gates, the quantum circuit has been transformed into a classical statistical model, where the degrees of freedom are formed by permutation-valued spins $\sigma$, $\tau$ on a honeycomb lattice. The partition function $Z$ of this statistical model is obtained by summing the total weights of different spin configurations. We further trace over all $\tau$ spins to make each configuration weight positive definite.
@@
The statistical model is ferromagnetic by considering the weights of each downward triangle with specific spin configurations as $W^{0}(\sigma,\sigma;\sigma)\approx d^{0}$ and $W^{0}(\sigma^{\prime},\sigma;\sigma)\approx d^{-|(\sigma^{\prime})^{-1})\sigma|}<d^{0}$ in the large $d$ limit, where $|\sigma|$ is the number of transpositions required to arrive $\sigma$ from the identity permutation. Therefore, all the spins tend to align in the same direction to achieve the largest total weight. However, as discussed below, due to the particular boundary conditions, domain walls may appear with free energy of $\log(W^{0}(\sigma^{\prime},\sigma;\sigma))$ for unit length. It is also worth noting that the weight $W^{0}(\sigma,\sigma;\sigma^{\prime})=0$ due to unitary constraint Dong et al. (2021); Li et al. (2023), which further restricts the possible configurations of horizontal domain walls.
@@
where $C=\begin{pmatrix}1&2&...&n\\ 2&3&...&1\end{pmatrix}$ and $I=\begin{pmatrix}1&2&...&n\\ 1&2&...&n\end{pmatrix}$ are the cyclic and identity permutations in $S_{n}$ group applied on each qudit, respectively. With the help of the replica trick Nishimori (2001); Kardar (2007), we can overcome the difficulty of averaging outside the logarithmic function
@@
    The least impacted factor is the separation of the annealed and quenched averaged results for finite $d$, i.e. the difference between $-\ln\mathbb{E}_{\mathcal{U}}Z$ and $-\mathbb{E}_{\mathcal{U}}\ln Z$. This difference can be properly addressed by considering the effective interaction between replicas and the term is in the order of $1/(d^{8}\ln d)$ as revealed in Zhou and Nahum (2019). Therefore, this difference can be safely ignored for most needs even with $d=2$. In the following, we only consider $k=1$ replica as we are satisfied with annealed average results.
@@
$\mathbb{E}_{2}(v)=\log_{d}\frac{d^{2}+1}{d}+\frac{1+v}{2}\log_{d}\frac{1+v}{2}+\frac{1-v}{2}\log_{d}\frac{1-v}{2}.$ (25)
@@
We now focus on some special limits of the entanglement membrane picture for a better understanding. Firstly, we consider the problem of half-chain entanglement growth in an infinitely long random quantum circuit. This quantity can map to the ferromagnetic statistical model with a top boundary condition of half chain in $\mathbb{C}$ and half chain in $\mathbb{I}$. Since the space extent is infinite, the dominant domain wall must be vertical. Noticing the bottom boundary is free due to the product state input, the vertical domain wall with velocity $v(t)=v$ gives the free energy as $E_{2}(v)T$. According to Eq. (25), the minimum value of $\mathbb{E}$ is achieved when $v=0$, we have $v_{E}=\mathbb{E}(0)=\log_{d}\frac{d^{2}+1}{2d}\approx 0.322$. Note that the definition of time unit in our work is twice as Zhou and Nahum (2019), i.e. our $v_{E}=0.644$. Further analysis including the difference between annealed and quenched averages gives a more accurate result as in Eq. (6) in Zhou and Nahum (2019), the speed of Rényi-2 entropy in random Haar circuit is given by the field theory up to $O(\frac{1}{d^{10}\ln d})$ as $v_{E}\approx 0.643$. We can confirm that the difference between annealed and quenched average is small as discussed in factor (a) above. The entanglement membrane picture is summarized in Fig. 3(a)(b).
@@
We now consider another limit, focusing on the entanglement entropy of the finite interval $L_{A}$. In late times, the entanglement is saturated, the statistical model has a top boundary condition with $L_{A}$ sites having $\mathbb{C}$ configuration and others having $\mathbb{I}$ configurations. Since now the time direction is sufficiently long, the vertical domain wall is not favorable anymore. The favored domain wall would be a downward triangle with velocity $v$ on both sides near the top boundary as shown in Fig. 3(c)(d). The free energy is then $\mathbb{E}_{2}(v)\frac{L_{A}}{v}$. To minimize this objective, we set the derivative zero: $v\mathbb{E}_{2}^{\prime}(v)=\mathbb{E}_{2}(v)$, this gives a velocity $v_{B}$, which corresponds to the operator spreading speed, satisfying the assumption that $\mathbb{E}_{2}(v_{B})=v_{B}$ and $\mathbb{E}_{2}^{\prime}(v_{B})=1$ Jonay et al. (2018). In random quantum circuit case, $v_{B}=\frac{d^{2}-1}{d^{2}+1}$. Considering the time unit definition is our work, we have $v_{B}=1.2$ in $d=2$ qubit system. Plugging the speed $v_{B}$ into the free energy, we have $S_{2}=L_{A}$. This result is particularly intriguing, as it contrasts with the vertical domain wall; the free energy of the triangular domain wall near the top boundary remains identical for the large $d$ limit and finite $d$ case. Therefore, in the following domain wall pictures, we simply use $v=1$ domain wall in the visualization, which represents the free energy contribution in terms of the spatial length $L_{A}$ both in the large $d$ and finite $d$ case.
III.2 Numerical results
@@
We now present the numerical results for different circuit models and different encoding schemes using random Clifford circuits Aaronson and Gottesman (2004), along with quantitative predictions based on the analytical framework presented above. We demonstrate that the numerical simulations align well with the analytical predictions.
III.2.1 Random circuit with one-to-one encoding

### v3

@@ 2.5.2 Relation with OTOC
The SIC proposed in this work has several advantages over the commonly used probe OTOC for diagnosing information scrambling. From experimental perspectives, SIC only requires one call on the evolution operator $U$, thereby eliminating the need for an extra invocation on reversed time evolution $U^{\dagger}$, which is typically required for measuring OTOCs [43]. (See Appendix D for detailed comparisons of experimental protocols). While OTOCs of local operators are frequently utilized to investigate when and how local information could reach distant locations, they in general may not resolve the detailed spatial distribution of conserved or slowly scrambling information within the system at late times. Consequently, they are not straightforward to differentiate phases without information scrambling. On the contrary, SIC provides a more comprehensive insight into the global structure of information propagation, evolution, and distribution and gives a finer classification on non-chaotic phases. Notably, this research is the first to unveil the steady-state information structures across a variety of dynamical phases with the lens of SIC.
2.5.3 Relation with tripartite mutual information
Tripartite mutual information has been proposed to explore the intrinsic information scrambling capacity of unitary evolutions [26]. In settings for TMI, all system qubits are entangled with external ancilla qubits, as already discussed in Sec. 2.5.2.
Indeed, if $\rho_{\bar{E}}$ defined in Eq. (7) is $I$ instead of a pure state $|\psi_{\bar{E}}\rangle\langle\psi_{\bar{E}}|$, our setup can be reduced to the case in [26], where TMI is given as
@@ 2.5.3 Relation with tripartite mutual information
From the experimental perspective, the fully mixed initial state preparation for standard TMI protocols is resource-intensive, typically demanding $2L$ qubits ($L$ system qubits entangled with $L$ reference qubits). Conversely, the SIC with one-to-one encoding scheme is considerably more efficient, requiring only a single reference qubit and a total of $L+1$ qubits.
2.5.4 Relation with Holevo information
Another closely related information metric is Holevo information [44]. The quantity provides a lower bound on the classical information transmission capacity for a quantum channel, analogous to how quantum coherent information bounds the quantum information transmission capacity. The definition of Holevo information, in terms of the effective subsystem channel introduced in this work, is
@@ 2.5.4 Relation with Holevo information
Accordingly, under the one-to-one encoding scheme, the first term of Holevo information is reduced to $S_{A}$ in the language of our main settings with reference qubits $R$. The second term of Holevo information is straightforward as the steady-state entanglement for subsystem $A$ averaged over different initial product states. According to the statistical model mapping introduced in Sec. 3.1, the second term differs from $S_{A}$ term in our main settings by the free bottom boundary condition for $E$. For comparison, the bottom boundary condition for $E$ for $S_{A}$ and $S_{AR}$ are $\mathbb{I}$ and $\mathbb{C}$, respectively (Notations on the effective statistical model are introduced in Sec. 3.1). Holevo information ranges from 0 to 1, with the two limits implying complete information loss and complete information protection, respectively.
@@
In the subsequent sections, we present our findings for the random circuit systems and Hamiltonian quench dynamics. The numerical data were obtained using state-of-the-art computational tools, including the Julia package QuantumClifford.jl for large-scale Clifford circuit simulations, the Python package TensorCircuit-NG [45] for scalable free fermion simulations, and the Python package QuSpin [46, 47] for simulations involving interacting fermions. The code implementation can be found at [48].
@@
3 Results for random circuit dynamics
Random quantum circuits [35] emerge as a promising and powerful platform for exploring a variety of non-equilibrium phenomena, particularly from information perspectives including entanglement growth [20, 49] and information scrambling characterized by OTOC and other metrics [50, 51, 52, 53, 54]. Recently, the noise-resilience phase in monitored random circuits that can protect information from mid-circuit measurements also attracted a lot of interest [36, 37, 38, 55, 56, 57, 58, 10, 39, 59, 11, 12, 60, 61, 62, 63]. We remark that mid-circuit measurement is the only ingredient we consider in this work that leads to open system evolution and information leakage, i.e. $I(|A|=L:R)(t)<2$. It is an interesting future direction to explore possibilities and properties of other gadgets leading to information leakage in information dynamics, which include quantum noises [64, 65, 66, 67, 68, 69] or construction with explicit ancillary environments [70].
To investigate the dynamical behavior in random circuits with the lens of SIC, we consider systems with qubits $d=2$ numerically, and in the theoretical analysis, we consider systems composed of $L$ $d$-qudits with both large $d$ limit and finite $d$ corrections ensuring a comprehensive understanding of the dynamics.
3.1 Analytical framework
In this subsection, we provide a concise review of the analytical method that establishes the correspondence between random quantum circuits and effective classical spin statistical models, with further details elaborated in Appendix C. Within this framework, the entanglement entropy is interpreted as a so-called entanglement membrane in space-time [71, 72, 73], i.e. minimal free energy domain walls between different permutation valued spin in one dimension case. For an in-depth discussion of the mapping and the field theory, we direct the reader to [74, 75, 76, 77, 64, 66].
The quantum circuit under investigation is composed of random two-qudit unitary gates arranged in a brickwall layout as shown in Fig 1(a). In this work, we adopt the convention by taking the unit time $\Delta t$ = 1 for an even-odd layer of two-qudit gates. Note that the convention gives the lightcone speed $v=2$, which may be different from some literature where $\Delta t=1$ refers to one even or one odd layer. After $T$ layers of gates, the output density matrix $\rho$ is
@@ 3.1 Analytical framework
where $S_{r}$ is the permutation group of degree $r$, $d$ is the local Hilbert space dimension of qudit, and $\text{Wg}_{d^{2}}^{(r)}$ is the Weingarten function [75, 78]. By independently averaging over all two-qudit gates, the quantum circuit has been transformed into a classical statistical model, where the degrees of freedom are formed by permutation-valued spins $\sigma$, $\tau$ on a honeycomb lattice. The partition function $Z$ of this statistical model is obtained by summing the total weights of different spin configurations. We further trace over all $\tau$ spins to make each configuration weight positive definite.
@@
The statistical model is ferromagnetic by considering the weights of each downward triangle with specific spin configurations as $W^{0}(\sigma,\sigma;\sigma)\approx d^{0}$ and $W^{0}(\sigma^{\prime},\sigma;\sigma)\approx d^{-|(\sigma^{\prime})^{-1})\sigma|}<d^{0}$ in the large $d$ limit, where $|\sigma|$ is the number of transpositions required to arrive $\sigma$ from the identity permutation. Therefore, all the spins tend to align in the same direction to achieve the largest total weight. However, as discussed below, due to the particular boundary conditions, domain walls may appear with free energy of $\log(W^{0}(\sigma^{\prime},\sigma;\sigma))$ for unit length. It is also worth noting that the weight $W^{0}(\sigma,\sigma;\sigma^{\prime})=0$ due to unitary constraint [79, 80], which further restricts the possible configurations of horizontal domain walls.
@@
where $C=\begin{pmatrix}1&2&...&n\\ 2&3&...&1\end{pmatrix}$ and $I=\begin{pmatrix}1&2&...&n\\ 1&2&...&n\end{pmatrix}$ are the cyclic and identity permutations in $S_{n}$ group applied on each qudit, respectively. With the help of the replica trick [81, 82], we can overcome the difficulty of averaging outside the logarithmic function
@@
    The least impacted factor is the separation of the annealed and quenched averaged results for finite $d$, i.e. the difference between $-\ln\mathbb{E}_{\mathcal{U}}Z$ and $-\mathbb{E}_{\mathcal{U}}\ln Z$. This difference can be properly addressed by considering the effective interaction between replicas and the term is in the order of $1/(d^{8}\ln d)$ as revealed in [75]. Therefore, this difference can be safely ignored for most needs even with $d=2$. In the following, we only consider $k=1$ replica as we are satisfied with annealed average results.
@@
$\mathbb{E}_{2}(v)=\log_{d}\frac{d^{2}+1}{d}+\frac{1+v}{2}\log_{d}\frac{1+v}{2}$
$+\frac{1-v}{2}\log_{d}\frac{1-v}{2}.$ (23)
@@
We now focus on some special limits of the entanglement membrane picture for a better understanding. Firstly, we consider the problem of half-chain entanglement growth in an infinitely long random quantum circuit. This quantity can map to the ferromagnetic statistical model with a top boundary condition of half chain in $\mathbb{C}$ and half chain in $\mathbb{I}$. Since the space extent is infinite, the dominant domain wall must be vertical. Noticing the bottom boundary is free due to the product state input, the vertical domain wall with velocity $v(t)=v$ gives the free energy as $E_{2}(v)T$. According to Eq. (23), the minimum value of $\mathbb{E}$ is achieved when $v=0$, we have $v_{E}=\mathbb{E}(0)=\log_{d}\frac{d^{2}+1}{2d}\approx 0.322$. Note that the definition of time unit in our work is twice as [75], i.e. our $v_{E}=0.644$. Further analysis including the difference between annealed and quenched averages gives a more accurate result as in Eq. (6) in [75], the speed of Rényi-2 entropy in random Haar circuit is given by the field theory up to $O(\frac{1}{d^{10}\ln d})$ as $v_{E}\approx 0.643$. We can confirm that the difference between annealed and quenched average is small as discussed in factor (a) above. The entanglement membrane picture is summarized in Fig. 3(a)(b).
@@
We now consider another limit, focusing on the entanglement entropy of the finite interval $L_{A}$. In late times, the entanglement is saturated, the statistical model has a top boundary condition with $L_{A}$ sites having $\mathbb{C}$ configuration and others having $\mathbb{I}$ configurations. Since now the time direction is sufficiently long, the vertical domain wall is not favorable anymore. The favored domain wall would be a downward triangle with velocity $v$ on both sides near the top boundary as shown in Fig. 3(c)(d). The free energy is then $\mathbb{E}_{2}(v)\frac{L_{A}}{v}$. To minimize this objective, we set the derivative zero: $v\mathbb{E}_{2}^{\prime}(v)=\mathbb{E}_{2}(v)$, this gives a velocity $v_{B}$, which corresponds to the operator spreading speed, satisfying the assumption that $\mathbb{E}_{2}(v_{B})=v_{B}$ and $\mathbb{E}_{2}^{\prime}(v_{B})=1$ [71]. In random quantum circuit case, $v_{B}=\frac{d^{2}-1}{d^{2}+1}$. Considering the time unit definition is our work, we have $v_{B}=1.2$ in $d=2$ qubit system. Plugging the speed $v_{B}$ into the free energy, we have $S_{2}=L_{A}$. This result is particularly intriguing, as it contrasts with the vertical domain wall; the free energy of the triangular domain wall near the top boundary remains identical for the large $d$ limit and finite $d$ case. Therefore, in the following domain wall pictures, we simply use $v=1$ domain wall in the visualization, which represents the free energy contribution in terms of the spatial length $L_{A}$ both in the large $d$ and finite $d$ case.
3.2 Numerical results
@@
We now present the numerical results for different circuit models and different encoding schemes using random Clifford circuits [83], along with quantitative predictions based on the analytical framework presented above. We demonstrate that the numerical simulations align well with the analytical predictions.
3.2.1 Random circuit with one-to-one encoding